OpenStudy (nuccioreggie):

Consider the following problem: Florence has 8 nieces and she wants to divide her collection of dolls among them evenly. If each niece receives thirty-two dolls, how many dolls were in Florence's collection? If d represents the number of dolls in Florence's collection, which of the following equations can be used to model the problem? 8 times d equals thirty-two thirty-two times d equals 8 thirty-two divided by 8 equals d d divided by 8 equals thirty-two

OpenStudy (nuccioreggie):

@mathmath333

OpenStudy (mathmath333):

\(\large\tt \begin{align} \color{red}{\text{there are total 8 nieces}\\~\\ 1~niece\implies 32 dolles\\~\\ \text{d dolls were in Florence's collection}\\ \implies 8\times 32~~ dolles\\~\\ \color{blue}{\dfrac{d}{8}=32}}\end{align}\)

OpenStudy (nuccioreggie):

so the answer is d @mathmath333

OpenStudy (mathmath333):

what do u think